# Global Energy Minimization for Simplex Mesh Optimization: A Radius Ratio Approach to Sliver Elimination

### v2

@@ 2 Radius Ratio Energy Function
$\nabla_{\bm{x}_{0}}\mu_{n}~=\frac{pq}{16|\tau_{n}|^{2}}(\frac{\nabla_{\bm{x}_{0}}p}{p}+\frac{\nabla_{\bm{x}_{0}}q}{q}-\frac{2}{|\tau_{n}|}\nabla_{\bm{x}_{0}}|\tau_{n}|)$
@@
$c=\frac{1}{|\tau_{n}|},c_{0}=\frac{1}{pl_{0}}+\frac{1}{l_{0}^{2}},c_{1}=~\frac{1}{pl_{1}}+\frac{1}{l_{1}^{2}},c_{2}=\frac{1}{pl_{2}}+\frac{1}{l_{2}^{2}},$
@@
$\nabla_{\bm{V}_{n}}\mu_{n}~=~\begin{bmatrix}\bm{A}_{n}~&~\bm{B}_{n}\\ -\bm{B}_{n}~&~\bm{A}_{n}\end{bmatrix}~\bm{V}_{n}$
@@
$\bm{A}_{n}=\mu_{n}~\begin{pmatrix}c_{12}~&~-c_{2}~&~-c_{1}~\\ -c_{2}~&~c_{20}~&~-c_{0}~\\ -c_{1}~&~-c_{0}~&~c_{01}\end{pmatrix},\bm{B}_{n}~=~\mu_{n}\begin{pmatrix}0~&~-c~&~c~\\ c~&~0~&~-c~\\ -c~&~c~&~0\end{pmatrix}$
@@
$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}~&~\bm{B}~\\ -\bm{B}~&~\bm{A}~\\ \end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\end{bmatrix}=\bm{G}_{F}\bm{V}$
@@
$\nabla_{\bm{x}_{0}}|\bm{d}_{0}|=$ $\frac{1}{|\bm{d}_{0}|}\{2[\bm{d}_{0}\cdot(\bm{v}_{20}\times\bm{v}_{30})\bm{v}_{10}+\bm{d}_{0}\cdot(\bm{v}_{30}\times\bm{v}_{10})\bm{v}_{20}+\bm{d}_{0}\cdot(\bm{v}_{10}\times\bm{v}_{20})\bm{v}_{30}]~$
@@
$[\bm{d}_{0}]_{\times}=\bm{D}_{0}=\begin{bmatrix}0~&~-d_{2}~&~d_{1}~\\ d_{2}~&~0~&~-d_{0}~\\ -d_{1}~&~d_{0}~&~0\end{bmatrix}$
@@
$\nabla_{\bm{V}_{n}}|\bm{d}_{0}|~=~\frac{1}{12|\tau_{n}|R_{n}}\begin{bmatrix}\bm{M}~&~-d_{2}\bm{K}~&~d_{1}\bm{K}~\\ d_{2}\bm{K}~&~\bm{M}~&~-d_{0}\bm{K}~\\ -d_{1}\bm{K}~&~d_{0}\bm{K}~&~\bm{M}\end{bmatrix}~\bm{V}_{n}$
@@
$\bm{M}=\begin{bmatrix}2c~&~-2c_{23}~&~-2c_{31}~&~-2c_{12}~\\ -2c_{23}~&~2c_{23}~&~0~&~0~\\ -2c_{31}~&~0~&~2c_{31}~&~0~\\ -2c_{12}~&~0~&~0~&~2c_{12}\end{bmatrix},\bm{K}=\begin{bmatrix}0~&~-k_{23}~&~-k_{31}~&~-k_{12}~\\ k_{23}~&~0~&~-\bm{v}_{30}^{2}~&~\bm{v}_{20}^{2}~\\ k_{31}~&~\bm{v}_{30}^{2}~&~0~&~-\bm{v}_{10}^{2}~\\ k_{12}~&~-\bm{v}_{20}^{2}~&~\bm{v}_{10}^{2}~&~0\end{bmatrix}$
@@
$p_{0}$ $=~\frac{\bm{v}_{31}^{2}}{4s_{2}}+\frac{\bm{v}_{21}^{2}}{4s_{3}}+\frac{\bm{v}_{32}^{2}}{4s_{1}},\quad~p_{1}~=~\frac{\bm{v}_{32}^{2}}{4s_{0}}+\frac{\bm{v}_{02}^{2}}{4s_{3}}+\frac{\bm{v}_{30}^{2}}{4s_{2}},\quad$
$~p_{2}$ $=~\frac{\bm{v}_{30}^{2}}{4s_{1}}+\frac{\bm{v}_{10}^{2}}{4s_{3}}+\frac{\bm{v}_{31}^{2}}{4s_{0}},\quad p_{3}~=~\frac{\bm{v}_{10}^{2}}{4s_{2}}+\frac{\bm{v}_{20}^{2}}{4s_{1}}+\frac{\bm{v}_{21}^{2}}{4s_{0}},\quad$
@@
$\nabla_{\bm{V}_{n}}\mu_{n}~=~\begin{bmatrix}\bm{S}~&~~&~~\\ ~&~\bm{S}~&~~\\ ~&~~&~\bm{S}\end{bmatrix}~\bm{V}_{n}$
@@
$\bm{S}~=~\begin{bmatrix}p_{0}&q_{01}&q_{02}&q_{03}\\ q_{01}&p_{1}&~q_{12}&~q_{13}\\ q_{02}&~q_{12}&p_{2}&q_{23}\\ q_{03}&q_{13}&q_{23}&p_{3}\end{bmatrix}$
@@
$\bm{C}_{0}=\begin{bmatrix}0~&~x_{2}~&~x_{3}~&~x_{1}~\\ x_{3}~&~0~&~x_{0}~&~x_{2}~\\ x_{1}~&~x_{3}~&~0~&~x_{0}~\\ x_{2}~&~x_{0}~&~x_{1}~&0\end{bmatrix},\bm{C}_{1}=\begin{bmatrix}0~&~y_{2}~&~y_{3}~&~y_{1}~\\ y_{3}~&~0~&~y_{0}~&~y_{2}~\\ y_{1}~&~y_{3}~&~0~&~y_{0}~\\ y_{2}~&~y_{0}~&~y_{1}~&0\end{bmatrix},\bm{C}_{2}=\begin{bmatrix}0~&~z_{2}~&~z_{3}~&~z_{1}~\\ z_{3}~&~0~&~z_{0}~&~z_{2}~\\ z_{1}~&~z_{3}~&~0~&~z_{i}~\\ z_{2}~&~z_{0}~&~z_{1}~&0\end{bmatrix}$
@@
$\nabla_{\bm{V}_{n}}|\tau_{n}|=\frac{1}{6}\begin{bmatrix}0~&~-\bm{C}_{2}~&~\bm{C}_{1}~\\ \bm{C}_{2}~&~0~&~-\bm{C}_{0}~\\ -\bm{C}_{1}~&~\bm{C}_{0}~&~0\end{bmatrix}\bm{V}_{n}=\bm{C}\bm{V}_{n}$
@@
$f~=W_{ik}^{j}~x_{i}~x_{j}x_{k}$
So we have $\nabla_{x_{k}}f=U_{ijk}x_{i}x_{j}=W_{ik}^{j}x_{i}x_{j}$, let $C_{ki}~=~U_{ijk}x_{j},~E_{ki}~=~W_{ik}^{j}x_{j}$, so $\bm{E}~=~\frac{1}{2}(\bm{C}+\bm{C}^{T})$, then we can express $\nabla f$ as
@@
$\frac{1}{2}(\bm{C}+\bm{C}^{T})=\frac{1}{12}\begin{bmatrix}0~&~\bm{C}_{2}^{T}-\bm{C}_{2}~&~\bm{C}_{1}^{T}-\bm{C}_{1}~\\ \bm{C}_{2}-\bm{C}_{2}^{T}~&~0~&~\bm{C}_{0}^{T}-\bm{C}_{0}~\\ \bm{C}_{1}^{T}-\bm{C}_{1}~&~\bm{C}_{0}^{T}-\bm{C}_{0}~&~0\end{bmatrix}$
@@
$\nabla_{\bm{V}_{n}}\mu_{n}=\mu_{n}\begin{bmatrix}\bm{A}_{n}~&~\bm{B}_{0n}~&~\bm{B}_{1n}~\\ -\bm{B}_{2n}~&~\bm{A}_{n}~&~\bm{B}_{0n}~\\ -\bm{B}_{1n}~&~-\bm{B}_{0n}~&~\bm{A}_{n}\end{bmatrix}\bm{V}_{n}$
@@
$\bm{A}_{n}=\frac{1}{12|\tau_{n}|R_{n}}\bm{M}+\frac{1}{s}\bm{S},~\bm{B}_{0n}=\frac{-d_{0}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{0}^{T}-\bm{C}_{0}),$
@@
$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}~&~\bm{B}_{2}~&~\bm{B}_{1}~\\ -\bm{B}_{2}~&~\bm{A}~&~\bm{B}_{0}~\\ -\bm{B}_{1}~&~-\bm{B}_{0}~&~\bm{A}\end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\\ \bm{Z}\end{bmatrix}=\bm{G}_{F}\bm{V}~$
@@ 3 Energy Optimization
$\bm{M}=\begin{bmatrix}2c~&~-2c_{23}~&~-2c_{31}~&~-2c_{12}~\\ -2c_{23}~&~2c_{23}~&~0~&~0~\\ -2c_{31}~&~0~&~2c_{31}~&~0~\\ -2c_{12}~&~0~&~0~&~2c_{12}\end{bmatrix}$
@@
$\bm{M}^{abs}=\begin{bmatrix}2|c_{12}|+|c_{23}|+|c_{31}|~&~-2|c_{23}|~&~-2|c_{31}|~&~-2|c_{12}|~\\ -2|c_{23}|~&~2|c_{23}|~&~0~&~0~\\ -2|c_{31}|~&~0~&~2|c_{31}|~&~0~\\ -2|c_{12}|~&~0~&~0~&~2|c_{12}|\end{bmatrix}$

### v3

@@ 2 Radius Ratio Energy Function
$\nabla_{\bm{x}_{0}}\mu_{n}\nobreakspace=\frac{pq}{16|\tau_{n}|^{2}}(\frac{\nabla_{\bm{x}_{0}}p}{p}+\frac{\nabla_{\bm{x}_{0}}q}{q}-\frac{2}{|\tau_{n}|}\nabla_{\bm{x}_{0}}|\tau_{n}|)$
@@
$c=\frac{1}{|\tau_{n}|},c_{0}=\frac{1}{pl_{0}}+\frac{1}{l_{0}^{2}},c_{1}=\nobreakspace\frac{1}{pl_{1}}+\frac{1}{l_{1}^{2}},c_{2}=\frac{1}{pl_{2}}+\frac{1}{l_{2}^{2}},$
@@
$\nabla_{\bm{V}_{n}}\mu_{n}\nobreakspace=\nobreakspace\begin{bmatrix}\bm{A}_{n}\nobreakspace&\nobreakspace\bm{B}_{n}\\ -\bm{B}_{n}\nobreakspace&\nobreakspace\bm{A}_{n}\end{bmatrix}\nobreakspace\bm{V}_{n}$
@@
$\bm{A}_{n}=\mu_{n}\nobreakspace\begin{pmatrix}c_{12}\nobreakspace&\nobreakspace-c_{2}\nobreakspace&\nobreakspace-c_{1}\nobreakspace\\ -c_{2}\nobreakspace&\nobreakspace c_{20}\nobreakspace&\nobreakspace-c_{0}\nobreakspace\\ -c_{1}\nobreakspace&\nobreakspace-c_{0}\nobreakspace&\nobreakspace c_{01}\end{pmatrix},\bm{B}_{n}\nobreakspace=\nobreakspace\mu_{n}\begin{pmatrix}0\nobreakspace&\nobreakspace-c\nobreakspace&\nobreakspace c\nobreakspace\\ c\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-c\nobreakspace\\ -c\nobreakspace&\nobreakspace c\nobreakspace&\nobreakspace 0\end{pmatrix}$
@@
$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}\nobreakspace&\nobreakspace\bm{B}\nobreakspace\\ -\bm{B}\nobreakspace&\nobreakspace\bm{A}\nobreakspace\\ \end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\end{bmatrix}=\bm{G}_{F}\bm{V}$
@@
$\nabla_{\bm{x}_{0}}|\bm{d}_{0}|=$ $\frac{1}{|\bm{d}_{0}|}\{2[\bm{d}_{0}\cdot(\bm{v}_{20}\times\bm{v}_{30})\bm{v}_{10}+\bm{d}_{0}\cdot(\bm{v}_{30}\times\bm{v}_{10})\bm{v}_{20}+\bm{d}_{0}\cdot(\bm{v}_{10}\times\bm{v}_{20})\bm{v}_{30}]\nobreakspace$
@@
$[\bm{d}_{0}]_{\times}=\bm{D}_{0}=\begin{bmatrix}0\nobreakspace&\nobreakspace-d_{2}\nobreakspace&\nobreakspace d_{1}\nobreakspace\\ d_{2}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-d_{0}\nobreakspace\\ -d_{1}\nobreakspace&\nobreakspace d_{0}\nobreakspace&\nobreakspace 0\end{bmatrix}$
@@
$\nabla_{\bm{V}_{n}}|\bm{d}_{0}|\nobreakspace=\nobreakspace\frac{1}{12|\tau_{n}|R_{n}}\begin{bmatrix}\bm{M}\nobreakspace&\nobreakspace-d_{2}\bm{K}\nobreakspace&\nobreakspace d_{1}\bm{K}\nobreakspace\\ d_{2}\bm{K}\nobreakspace&\nobreakspace\bm{M}\nobreakspace&\nobreakspace-d_{0}\bm{K}\nobreakspace\\ -d_{1}\bm{K}\nobreakspace&\nobreakspace d_{0}\bm{K}\nobreakspace&\nobreakspace\bm{M}\end{bmatrix}\nobreakspace\bm{V}_{n}$
@@
$\bm{M}=\begin{bmatrix}2c\nobreakspace&\nobreakspace-2c_{23}\nobreakspace&\nobreakspace-2c_{31}\nobreakspace&\nobreakspace-2c_{12}\nobreakspace\\ -2c_{23}\nobreakspace&\nobreakspace 2c_{23}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace\\ -2c_{31}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2c_{31}\nobreakspace&\nobreakspace 0\nobreakspace\\ -2c_{12}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2c_{12}\end{bmatrix},\bm{K}=\begin{bmatrix}0\nobreakspace&\nobreakspace-k_{23}\nobreakspace&\nobreakspace-k_{31}\nobreakspace&\nobreakspace-k_{12}\nobreakspace\\ k_{23}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-\bm{v}_{30}^{2}\nobreakspace&\nobreakspace\bm{v}_{20}^{2}\nobreakspace\\ k_{31}\nobreakspace&\nobreakspace\bm{v}_{30}^{2}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-\bm{v}_{10}^{2}\nobreakspace\\ k_{12}\nobreakspace&\nobreakspace-\bm{v}_{20}^{2}\nobreakspace&\nobreakspace\bm{v}_{10}^{2}\nobreakspace&\nobreakspace 0\end{bmatrix}$
@@
$p_{0}$ $=\nobreakspace\frac{\bm{v}_{31}^{2}}{4s_{2}}+\frac{\bm{v}_{21}^{2}}{4s_{3}}+\frac{\bm{v}_{32}^{2}}{4s_{1}},\quad\nobreakspace p_{1}\nobreakspace=\nobreakspace\frac{\bm{v}_{32}^{2}}{4s_{0}}+\frac{\bm{v}_{02}^{2}}{4s_{3}}+\frac{\bm{v}_{30}^{2}}{4s_{2}},\quad$
$\nobreakspace p_{2}$ $=\nobreakspace\frac{\bm{v}_{30}^{2}}{4s_{1}}+\frac{\bm{v}_{10}^{2}}{4s_{3}}+\frac{\bm{v}_{31}^{2}}{4s_{0}},\quad p_{3}\nobreakspace=\nobreakspace\frac{\bm{v}_{10}^{2}}{4s_{2}}+\frac{\bm{v}_{20}^{2}}{4s_{1}}+\frac{\bm{v}_{21}^{2}}{4s_{0}},\quad$
@@
$\nabla_{\bm{V}_{n}}\mu_{n}\nobreakspace=\nobreakspace\begin{bmatrix}\bm{S}\nobreakspace&\nobreakspace\nobreakspace&\nobreakspace\nobreakspace\\ \nobreakspace&\nobreakspace\bm{S}\nobreakspace&\nobreakspace\nobreakspace\\ \nobreakspace&\nobreakspace\nobreakspace&\nobreakspace\bm{S}\end{bmatrix}\nobreakspace\bm{V}_{n}$
@@
$\bm{S}\nobreakspace=\nobreakspace\begin{bmatrix}p_{0}&q_{01}&q_{02}&q_{03}\\ q_{01}&p_{1}&\nobreakspace q_{12}&\nobreakspace q_{13}\\ q_{02}&\nobreakspace q_{12}&p_{2}&q_{23}\\ q_{03}&q_{13}&q_{23}&p_{3}\end{bmatrix}$
@@
$\bm{C}_{0}=\begin{bmatrix}0\nobreakspace&\nobreakspace x_{2}\nobreakspace&\nobreakspace x_{3}\nobreakspace&\nobreakspace x_{1}\nobreakspace\\ x_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace x_{0}\nobreakspace&\nobreakspace x_{2}\nobreakspace\\ x_{1}\nobreakspace&\nobreakspace x_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace x_{0}\nobreakspace\\ x_{2}\nobreakspace&\nobreakspace x_{0}\nobreakspace&\nobreakspace x_{1}\nobreakspace&0\end{bmatrix},\bm{C}_{1}=\begin{bmatrix}0\nobreakspace&\nobreakspace y_{2}\nobreakspace&\nobreakspace y_{3}\nobreakspace&\nobreakspace y_{1}\nobreakspace\\ y_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace y_{0}\nobreakspace&\nobreakspace y_{2}\nobreakspace\\ y_{1}\nobreakspace&\nobreakspace y_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace y_{0}\nobreakspace\\ y_{2}\nobreakspace&\nobreakspace y_{0}\nobreakspace&\nobreakspace y_{1}\nobreakspace&0\end{bmatrix},\bm{C}_{2}=\begin{bmatrix}0\nobreakspace&\nobreakspace z_{2}\nobreakspace&\nobreakspace z_{3}\nobreakspace&\nobreakspace z_{1}\nobreakspace\\ z_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace z_{0}\nobreakspace&\nobreakspace z_{2}\nobreakspace\\ z_{1}\nobreakspace&\nobreakspace z_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace z_{i}\nobreakspace\\ z_{2}\nobreakspace&\nobreakspace z_{0}\nobreakspace&\nobreakspace z_{1}\nobreakspace&0\end{bmatrix}$
@@
$\nabla_{\bm{V}_{n}}|\tau_{n}|=\frac{1}{6}\begin{bmatrix}0\nobreakspace&\nobreakspace-\bm{C}_{2}\nobreakspace&\nobreakspace\bm{C}_{1}\nobreakspace\\ \bm{C}_{2}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-\bm{C}_{0}\nobreakspace\\ -\bm{C}_{1}\nobreakspace&\nobreakspace\bm{C}_{0}\nobreakspace&\nobreakspace 0\end{bmatrix}\bm{V}_{n}=\bm{C}\bm{V}_{n}$
@@
$f\nobreakspace=W_{ik}^{j}\nobreakspace x_{i}\nobreakspace x_{j}x_{k}$
So we have $\nabla_{x_{k}}f=U_{ijk}x_{i}x_{j}=W_{ik}^{j}x_{i}x_{j}$, let $C_{ki}\nobreakspace=\nobreakspace U_{ijk}x_{j},\nobreakspace E_{ki}\nobreakspace=\nobreakspace W_{ik}^{j}x_{j}$, so $\bm{E}\nobreakspace=\nobreakspace\frac{1}{2}(\bm{C}+\bm{C}^{T})$, then we can express $\nabla f$ as
@@
$\frac{1}{2}(\bm{C}+\bm{C}^{T})=\frac{1}{12}\begin{bmatrix}0\nobreakspace&\nobreakspace\bm{C}_{2}^{T}-\bm{C}_{2}\nobreakspace&\nobreakspace\bm{C}_{1}^{T}-\bm{C}_{1}\nobreakspace\\ \bm{C}_{2}-\bm{C}_{2}^{T}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace\bm{C}_{0}^{T}-\bm{C}_{0}\nobreakspace\\ \bm{C}_{1}^{T}-\bm{C}_{1}\nobreakspace&\nobreakspace\bm{C}_{0}^{T}-\bm{C}_{0}\nobreakspace&\nobreakspace 0\end{bmatrix}$
@@
$\nabla_{\bm{V}_{n}}\mu_{n}=\mu_{n}\begin{bmatrix}\bm{A}_{n}\nobreakspace&\nobreakspace\bm{B}_{0n}\nobreakspace&\nobreakspace\bm{B}_{1n}\nobreakspace\\ -\bm{B}_{2n}\nobreakspace&\nobreakspace\bm{A}_{n}\nobreakspace&\nobreakspace\bm{B}_{0n}\nobreakspace\\ -\bm{B}_{1n}\nobreakspace&\nobreakspace-\bm{B}_{0n}\nobreakspace&\nobreakspace\bm{A}_{n}\end{bmatrix}\bm{V}_{n}$
@@
$\bm{A}_{n}=\frac{1}{12|\tau_{n}|R_{n}}\bm{M}+\frac{1}{s}\bm{S},\nobreakspace\bm{B}_{0n}=\frac{-d_{0}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{0}^{T}-\bm{C}_{0}),$
@@
$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}\nobreakspace&\nobreakspace\bm{B}_{2}\nobreakspace&\nobreakspace\bm{B}_{1}\nobreakspace\\ -\bm{B}_{2}\nobreakspace&\nobreakspace\bm{A}\nobreakspace&\nobreakspace\bm{B}_{0}\nobreakspace\\ -\bm{B}_{1}\nobreakspace&\nobreakspace-\bm{B}_{0}\nobreakspace&\nobreakspace\bm{A}\end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\\ \bm{Z}\end{bmatrix}=\bm{G}_{F}\bm{V}\nobreakspace$
@@ 3 Energy Optimization
$\bm{M}=\begin{bmatrix}2c\nobreakspace&\nobreakspace-2c_{23}\nobreakspace&\nobreakspace-2c_{31}\nobreakspace&\nobreakspace-2c_{12}\nobreakspace\\ -2c_{23}\nobreakspace&\nobreakspace 2c_{23}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace\\ -2c_{31}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2c_{31}\nobreakspace&\nobreakspace 0\nobreakspace\\ -2c_{12}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2c_{12}\end{bmatrix}$
@@
$\bm{M}^{abs}=\begin{bmatrix}2|c_{12}|+|c_{23}|+|c_{31}|\nobreakspace&\nobreakspace-2|c_{23}|\nobreakspace&\nobreakspace-2|c_{31}|\nobreakspace&\nobreakspace-2|c_{12}|\nobreakspace\\ -2|c_{23}|\nobreakspace&\nobreakspace 2|c_{23}|\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace\\ -2|c_{31}|\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2|c_{31}|\nobreakspace&\nobreakspace 0\nobreakspace\\ -2|c_{12}|\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2|c_{12}|\end{bmatrix}$